# Aligning LLMs to Ask Good Questions A Case Study in Clinical Reasoning

### v1

@@ E.1 Expert Annotations
Due to recruitment constraints, we collected at least one expert annotation per sample such that 23 samples were annotated by one expert, 138 samples annotated by two experts and 137 samples were annotated by more than two experts, resulting in a total of 724 annotations over 295 samples, excluding samples annotated during recruitment.
@@
Fleiss’ Kappa 0.59 Moderate 0.035
AC1 0.66 Substantial 0.033
Table 8: Inter-rater agreement on correct answers chosen for each generated questions.
@@ Results.
On 298 samples, excluding 4 test questions, all questions were considered plausible and correct answers generated by o1 showed 88.6% accuracy based on based on majority vote on 298 samples. Upon examining agreement on 295 samples, excluding 7 samples used during recruitment, we see moderate to substantial agreement (Wongpakaran et al., 2013; Wong et al., 2021) as shown in Table 8.
@@ Appendix F Expert Manual Evaluation
F.1 Recruitment
We used Prolific for recruitment and selected participants who indicated their healthcare employment role as Physician, Nurse, Pharmacist, and Paramedics in the US. Recruitment task included 7 questions from multiple choice annotation task as described in Appendix E.1 and expertise information questions. As the roles on Prolific are self-reported, we added four test questions to filter experts, annotated by two medical researchers. We varied the difficulty of each of the four questions and weighed the most difficult question as 2 points and rest as 1 point. We passed participants who received greater than 2 points. We asked participants for information indicating expertise such as current roles, number of years studying medicine, and number of years practicing medicine. From 40 participants who responded to our task, we passed a total of 12 participants, which included 4 Nurses, 2 Pharmacist, 2 Physician, and 1 Medical Assistant.
F.2 Ranking Annotation Task Setup
@@
We randomly selected 100 samples and asked participants to rank 5 different questions per sample from best to worst follow-up questions. The 5 questions were randomly selected out of 6 variations to reduce cognitive load and annotation time (Wang et al., 2022). The six question variations included one human-written question and five model generated questions under different experimental setup (Base Model, Alfa DPO Policy, Alfa DPO Coarse, Alfa PPO Policy, Alfa PPO Reward). At least 2, maximum of 3, human experts were randomly assigned for each sample. We paid participants 20 USD/hr.
@@
Figure 7: Expert preference ranking results showing pairwise win-rates.
@@
Figure 8: An example of majority annotated best question (PPO Reward) and worst question (DPO Coarse).
F.3 Results
@@
Similar to the results discussed in §6.4, human annotation win-rate shows prefrence towards Alfa PPO Reward, trained on fine-grained reward fusion as shown in Figure 7, which, as shown in qualitative example in Figure 8, demonstrates superior quality. Interestingly, while on automatic evaluation on medical diagnostic accuracy, Alfa DPO Policy outperformed other integration strategies, we observe that DPO trained with coarse attributes outperforms it on human annotations by a slight margin. This diverging outcome could indicate various factors, including stylistic preferences (Zhang et al., 2024b), that might influence annotation decisions which require further investigation.
Appendix G Prompts
G.1 Counterfactual Perturbations
@@
G.2 Automatic Synthetic Data Verification
@@
G.3 MediQ-AskDocs Task Construction

### v2

@@ C.2 Training Hyperparameters
Training times on single A100 GPU with batch size 256:
  • SFT: 4hr (3B model), 8.5hr (8B model)
  • DPO: 10hr (3B model), 42hr (8B model)
  • PPO: 48hr (3B model)
@@ E.1 Expert Annotations
Due to recruitment constraints, we collected three expert annotation per sample over 295 samples, excluding samples annotated during recruitment.
@@
Fleiss’ Kappa 0.58 Moderate 0.029
AC1 0.65 Substantial 0.027
Table 9: Inter-rater agreement on multiple-choice question correct answers.
@@ Results.
On 298 samples, excluding 4 test questions, all questions were considered plausible and correct answers generated by o1 showed 85.9% accuracy based on majority vote on 298 samples. Upon examining agreement on 295 samples, excluding 7 samples used during recruitment, we see moderate to substantial agreement (Wongpakaran et al., 2013; Wong et al., 2021) as shown in Table 9.
@@ Appendix F Expert Manual Evaluation
Expert manual evaluation was conducted by a panel of three research team members with extensive medical training and qualifications: two MD-PhD specialists with expertise in radiology, pathology, gastroenterology, and oncology, and one MD resident specializing in pathology. The members were instructed to rank the proposed questions given context (patient post and previous conversation—questions and answers between medical expert and patient—if any) in an online medical consultation setting similar to r/AskDocs.
*PO POI Win-rate
DPO Data 52
Policy 41
PPO Data 75
Reward 74
Policy 40
Human 47
Table 10: Win-rate of each model variation over baseline model for different attribute integration strategies based on majority vote.
*PO POI AC1 CI Lower CI Upper p-value PA PE
DPO Data 0.129 -0.090 0.349 0.245 0.520 0.449
Policy 0.240 0.047 0.434 0.016* 0.620 0.500
PPO Data 0.721 0.594 0.848 <0.001*** 0.780 0.211
Reward 0.680 0.543 0.817 <0.001*** 0.750 0.219
Policy 0.224 0.028 0.419 0.025* 0.610 0.498
*p<0.05, **p<0.01, ***p<0.001
Table 11: Gwet’s AC1 inter-rater reliability results between Human and LLM Judge on win-rate over baseline by model variation. PA denotes percent agreement and PE denotes expected percent agreement by chance.
F.1 Ranking Annotation Task Setup
@@
We randomly selected 100 samples and asked our panel to rank 7 different questions per sample from best to worst follow-up questions. The seven question variations included one human-written question and six model generated questions under different experimental setup (Base Model, Alfa DPO Data, Alfa DPO Policy, Alfa PPO Data, Alfa PPO Policy, Alfa PPO Reward). The questions were presented in a randomly shuffled order, and annotation platform was constructed to not allow ties.
@@
Figure 7: Expert preference ranking results showing pairwise win-rates. Models on y-axis are compared to the models on x-axis (e.g., Alfa PPO Data has 75% win-rate over Base Model).
@@
Figure 8: An example of majority annotated best question (PPO Reward) and worst question (DPO Policy).
F.2 Results
@@
Similar to the results discussed in §6.4, human annotation win-rate shows preference towards Alfa PPO Data and Alfa PPO Reward, trained on fine-grained reward fusion as shown in Table 10. Interestingly, while on automatic evaluation on medical diagnostic accuracy, Alfa DPO Policy outperformed other integration strategies, we observe that it only outperforms Alfa PPO Policy. This diverging outcome could indicate various factors, including stylistic preferences (Zhang et al., 2024b) that might influence annotation decisions which require further investigation. As noted in §6.4, our medical diagnostic accuracy evaluation relies on patient provided information; however, as shown in an example in Figure 8, the best quality question from Alfa PPO Reward is more open-ended and targets multiple aspects, which might not have been answered by the patient in our data.
Inter-rater Agreement
Medical experts had a substantial pairwise ranking agreement of .481.481 Gwet’s AC1 score. The experts also showed high agreement with LLM judge, especially for model variations with higher win-rate. Experts noted that questions with lower win-rate showed similar quality leading to random assignment of preferences as ties were not allowed on our platform, which could explain the lower agreement for model variations with lower win-rate.
Context
Patient: In Hospital Again…Need Real Life Dr. House 40m, white, non smoker casual drinker. I am currently in a hospital bed on a Heparin Drip. I have Vascular, Hematology, oncology, and even Gastro all stumped. Basically, I am under attack by blood clots but no one has a clue why. A year of this. Long story short - calf DVT 10/20. Found May Thurner, stent placed, clot resolved, moved on with life. Taken off thinners. July - spinal surgery. 3 weeks later, blood clots EVERYWHERE, spent a week in hospital had thrombectomy. Simple answer - surgery caused clots (even though ive had many surgeries before with no clots). Put on Xarelto. November - found more clots even on Xarelto. Do tPA lysis catheter puts me in hospital another week. Clears clogged stents, put on Lovenox. 2 weeks later - more clots. Just had another thrombectomy. I have tested negative for everything. Negative for all genetic factors, negative for rare clotting disorders. And yet, here I am. CT Scan of chest and pelvis no tumors. Lymph nodes swollen but s̈ubcentimeterïn the area where all the clotting keeps happening. Everyone is stumped. They won’t let me out of the hospital. I feel fine except for my leg which has been through alot. Only info to go off of - I have been anemic off and on, Monocyte % slighly high, folate very low, absolute immature gran count high. I am here in the hospital. I can ask anything. Give me your wildest ideas. Something to have them test. I want to go home but also desperate for a diagnosis. Thank you! PS: WOW- thank you all. Updates from what I’ve relayed off of this thread - hematology is now done with me unless something changes. They say they’ve tested for every clotting disorder and looked as much as they could for underlying cancers. They say definitively it is not APS. Update 2 (I don’t think anyone is still checking this but just in case) - Tumor marker test was negative but they did find a precancerous polyp from the colonoscopy. Doctor: Yup, was wondering for their clotting disorders if they tested for connective tissue disease and things like antiphospholipid, lupus anticoagulant etc Patient: Yes. They said i had a weak positive for APS. They dont think its that but said they were going to run some more tests but havent yet
Alfa-PPO-Data Have you had any of the following symptoms in the past 2 weeks before the clots occurred: pain or swelling of the chest, abdomen, or groin, difficulty breathing, coughing or fever, pale gums, yellowing of the skin or eyes, nosebleeds or bleeding gums?
Alfa-PPO-Reward Can you describe the anemia you’ve experienced? Was it constant, or did it only occur during periods of clotting, and if so, how often did it occur? Was it painful or associated with any other symptoms?
Alfa-PPO-Policy Did you ever get the LVT?
Alfa-DPO-Policy And have you had your thyroid checked?
Base So they ran the APS test but aren’t convinced it’s the cause. In that case, here are some wild ideas to consider: 1. **Autoimmune response to something else**: Despite having low folate, maybe the issue isn’t with folate deficiency but with an autoimmune response to something else. Could it
Alfa-DPO-Data That’s the problem - a weak positive is not conclusive, and they’re still awaiting more results or further testing to rule it out or confirm it. Okay, wild ideas time! Here are a few potential test ideas, not that they’re guaranteed to work, but might be considered: 1. **Cardiolip
Human Did you or your sister have any negative tests before being diagnosed with lupus / APS?
Table 12: Qualitative example of generated questions. The questions are ordered from best (Alfa-PPO-Data) to worst (Human) quality based on human annotation according to majority vote.
Appendix G Qualitative Analysis
Table 12 shows an example of a multi-turn conversation and generated follow-up questions ordered by ranking of best to worst quality as denoted by our team of medical experts. Alfa-PPO-Data ranked best out of all options and asks a comprehensive and targeted follow-up question. Moreover, the example shows that even while prompted to ask questions, and trained to do so, asking follow-up questions is a difficult task as shown by the response by base model and Alfa-DPO-Data.
Appendix H Prompts
H.1 Counterfactual Perturbations
@@
H.2 Automatic Synthetic Data Verification
@@
H.3 MediQ-AskDocs Task Construction